1. A mass of 7 kg stretches a spring 0.21 m from its equilibrium position. The mass is removed and another body of mass 0.50 kg is suspended from its equilibrium the source spring. If the spring is now stretched and released, what is its period of motion?
1
Expert's answer
2020-10-23T11:51:27-0400

Given:

"m_1=7\: kg"

"x_1=0.21 \:m"

"m_2=0.50\: kg"

"T-?"

The spring stiffness is equal

"k=\frac{m_1g}{x_1}=\rm\frac{7\times 9.8}{0.21}=326.7\: N/m"

The period of oscillation

"T=2\pi\sqrt{\frac{m_2}{k}}=6.28\sqrt{\frac{0.50}{326.7}}=0.25\:\rm s"